# SonifyAR: Context-Aware Sound Generation in Augmented Reality

### v1

@@ 3.2. Acquisition of AR Sound
While the above design space helps highlight the dimensions of AR sound and current authoring support, below we reflect on key challenges related to the limitations of sound acquisition methods in terms of handling these different dimensions.
For example, a mechanical noise of a virtual robot (virtuality) can be easily sourced from local or online sound databases. However, more specific sounds, like a virtual steel ball hitting a physical concrete wall surface or a virtual racecar jumping into a backyard pool (virtuality-reality) requires physical world understanding and material awareness. The space of possible sound effects here is near infinite. Thus, in such cases, generating sounds using text-to-sound models is preferable.
Conversely, certain highly recognizable sounds, like a dog barking, pose difficulties for current generative models, often resulting in outputs that are unsatisfactory and noisy. The complexity in user needs and trade-offs in different methods suggests that the combination of both generation and retrieval could serve as a practical solution.
@@ 3.3. Event Representation
To combine generation and retrieval methods, one challenge is how to condition this suite of different sound acquisition methods. While visual formats such as images (10096023, ) and videos (Zhou_2018_CVPR, ; ghose2022foleygan, ) have been widely used as inputs for sound generation models, they fall short in fully capturing AR context information, particularly user actions. To overcome this, we explore the use of text as a universal representation for sound acquisition inputs, which has the following benefits. First, text can sufficiently and precisely convey necessary context information and the specifics of sound-generating events (e.g.,“User slides a ceramic cup on a wooden surface”). Second, given that the AR experience operates within a hardware-software system with multi-modal sensors (e.g., camera, IMU, GPS), we can leverage this system to monitor and log events within the AR experience and summarize them into descriptive text. Third, recent advancements, as demonstrated by works like HuggingGPT (shen2023hugginggpt, ) which uses a LLM to control AI tasks, showcase the practicality of using LLMs as controllers to process text for sound generation.
@@ 4. SonifyAR Implementation
We introduce SonifyAR, a novel PbD sound authoring framework that uses context recognition and generative AI to create personalized, context-sensitive sounds for AR interactions (Figure 3). SonifyAR consists of three primary components : (1) Event Textualization: Every user action is recognized as an AR event, and the context of these events is transformed into textual descriptions. (2) Sound Acquisition: An LLM-controlled sound acquisition process that utilize four methods to produces sound effects based on the context. (3) User Interface: An interface that allows users to experience the AR scene and provides the capability to view, modify, and test sound effects for AR events.
@@ 4.1. Event Textualization
In SonifyAR, we utilize a PbD authoring framework to capture potential sound-producing AR events and compile the context information into text. When users interact with the AR space, SonifyAR captures AR events that can lead to sound feedback. Here we define an AR event as an user action and its subsequent result (e.g. when a user taps a virtual object to trigger its animation). For each AR event, we describe it with an event type (e.g. tapping an object ), action source (e.g. user or virtual object), and action target (e.g. virtual object or real-world plane). Information about the involved parties, such as virtual objects or real-world planes, will also be included in the context information. This collected event data is then transformed into text with a template.
@@
To support a broad range of AR sound interaction, we implemented six event types:
@@
    Collide: A virtual object collides with another or with a real-world surface. We implemented a specialized collision detection mechanism as introduced in subsection 4.4
@@
SonifyAR utilizes ARKit (apple-arkit-doc, )’s plane detection functionality (apple-arkit-plane-tracking, ) to assess the surrounding environment. Since most AR experiences are anchored to planes, the detected plane information serves as a crucial context. To enrich such information, we employ the Deep Material Segmentation model (upchurch_dense_2022, ) to segment the scene and identify the material of planes (Figure 3), which helps produce realistic sound effects when an AR event involves this plane. Currently, we support six materials: wood, carpet, concrete, paper, metal, and glass. Surfaces that do not fall into one of these six categories are labeled as unknown surface.
@@
The semantics of the virtual object is also crucial for sound generation. To ensure sound assets align with the virtual object’s material and its state, we collect a text description for all the virtual objects (e.g. ”This model is a toy robot made of metal.”) and their animations (e.g. ”A toy robot walks.”). These descriptions can be provided by the asset creator or, if necessary, we prompt the user to add relevant details. Any AR event involving the virtual object or its animation will incorporate these text descriptions.
@@
To aggregate the multi-source context information described above, we employ a simple text template: ”This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities]”.
In our template, Event Type is the type of event described by the type name, as listed above. Source refers to the subject of the event, which could be the user when the event is directly triggered by user, or virtual object when it interacts with real-world environment. Target Object is the object of the event, e.g. the plane that gets tapped on, or the animation that gets played. Additional Information on Involved Entities include details that further elucidate the source object, the triggered event, and the target object. Examples include material descriptions and animation details. We leave the corresponding entity field blank when events are missing a target object or additional information.
During user interactions, the system actively monitors and logs the events happening in AR space. The context information is fetched and plugged into the text template, crafting a coherent description that reflects the user’s interaction within the AR environment.
@@ 4.2. Sound Acquisition
SonifyAR utilizes the LLM to automatically retrieve or generate context-matching sound assets of an AR event. Inspired by HuggingGPT (shen2023hugginggpt, ) and Visual ChatGPT (wu2023visual, ), we utilize the LLM as a controller of multiple sound authoring methods. The LLM takes the text description of the event as input and replies with commands for multiple sound acquisition methods (Figure 5). At our current stage, we support four major sound authoring methods: local recommendation, online retrieval, text2sound generation, and text-guided sound transfer.
Local Recommendation
@@
The LLM can recommend sound assets stored in the local database based on semantics in the event description. Similar to other AR authoring tools, we collect a set of sound effects from Adobe Audition sound effects (adobe-audition, ), each labeled with a descriptive file name, like “Crash Aluminum Tray Bang” or “Liquid Mud Suction”. The entire list of sound file name is provided to LLM. When provided with the event context, the LLM recommends the top five most suitable local sound effects based on their file names. It then returns the selected sound effect in the format of method1recommend:FILENAME, where the FILENAME would be replaced with the actual filename. Upon receiving, SonifyAR parses the filename and adds the top corresponding sound as one of the sound options for the event. User can also long hold such option to reveal other results from the top5 recommendation.
Online Retrieval
@@
We also expand our retrieval capability with an online sound asset database called FreeSound (freesound, ). We use the FreeSound API, which returns a list of best matching sound effects based on a given query. The queries are condensed versions of full event descriptions generated by the LLM, returned in the format method2retrieval:PROMPT, with PROMPT replaced by the specific search query. The returned results are JSON strings containing information on all search results. We select the top five matches from the entire set of sound effects returned by the API, which are then downloaded and presented to the user as sound options.
Sound Generation
@@
We adopt a audio diffusion model AudioLDM (liu2023audioldm, ) to generate sound effects with text prompt. We ask the LLM to compress the event text description into a shortened generation prompt (returned in the format method3generation:PROMPT, where the PROMPT would be replaced with the generation prompt). Upon receiving such a command from the LLM, SonifyAR sends the prompt to the AudioLDM model, requesting text2sound generation. These newly generated sounds are then presented to user as sound options.
Sound Transfer
@@
Besides text2audio generation, AudioLDM is also capable of performing text-based sound style transfer. Specifically, for events like tapping, sliding, or colliding, instead of generating a new sound from scratch, SonifyAR uses a default sound effect and initiate a style transfer operation with a text prompt provided by the LLM. This approach allows the output sounds to match the length and rhythm of the input sounds, so that they can be well-timed with actions. Furthermore, when users wish to further modify any provided sound assets, SonifyAR offers text-based style transfer as a fine-tuning and customizing option.
All sound authoring processes listed above operate concurrently in the backend. This ensures uninterrupted AR experience.
@@ 4.3. User Interface
As a PbD authoring framework, the SonifyAR application initiate directly with an interactive AR experience (Figure 6 left). Users can freely interact with the scene, performing actions like moving virtual objects or interacting with the real-world surfaces. When an AR event is detected, a text label appears on the top-left screen to inform the user (Figure 6. A) and the sound acquisition component is activated to generate candidate sound effects for the detected events.
By clicking a button (Figure 6. D), users can prompt an editing interface (Figure 6. F) that overlays on the AR scene. The interface includes all unique AR events detected (Figure 6. G) and the automatically generated sound assets (Figure 6. H), enabling users to explore the detected events and their corresponding sound effects. Users can click on a sound effect to preview it and double click to select and activate it during the AR session. After confirming their choices, users can hide the authoring panel to resume the AR experience and test the selected sounds. Users can always access the editing interface to modify their choices.
Furthermore, users can long press a sound asset to call out a menu with a suite of exploratory options. For recommended or retrieved sounds, the menu includes an option to list all other sound assets in the top five recommendations. For all sound assets, the menu provides a “style transfer” and a “generate similar sounds” feature, enabling users to style transfer audio effects or to generate similar sounds based on the selected sound. When this feature is selected, users can type a simple text prompt to guide the sound generation process. This feature enables users to iteratively refine sound effects and explore new sound variations.
@@ 4.4. Technical Implementation
We developed SonifyAR as a mobile app with a user interface that enables AR interaction and sound authoring actions. We also implemented back-end servers for scene understanding and sound generation. The current SonifyAR demo is built with Apple’s ARKit with Swift 5.7 and Xcode 14, and runs as an iOS application on iPhones. Due to the enhanced performance of plane detection on devices with LiDAR (available in iPhone’s Pro lineup starting from iPhone Pro 12), we developed and tested the app on an iPhone 13 Pro Max running iOS 16. We implemented the collision detection between virtual objects and planes with ARKit’s physics simulation (apple-scenekit-physics-simulation, ) function. To enable collision detection for animated virtual object parts (e.g. the stomping of a robot foot on the plane), we bind colliders to the joints of the virtual object’s animated bones.
We use GPT-4 (openai2024gpt4, ) as the LLM in our system. We adopt the DMS model (upchurch_dense_2022, ) for material segmentation because of its detailed material labeling, fast processing time, and high accuracy. For sound generation, we use the AudioLDM model (liu2023audioldm, ) due to its state-of-the-art performance and its versatility in text2sound generation and text-guided sound2sound transfer. All models are hosted on a server equipped with an NVIDIA RTX 4080 GPU and CUDA 12.2.
@@ 5.1. Participants
We recruit eight participants (six male and two female, aged between 26 and 33) with varying experience in AR. We inquired about their prior experience in AR before the study. Out of the eight participants, six had prior AR experience, and three had experience in AR authoring tools.
@@ 5.2. Procedure
The usability study was conducted in a small meeting room and comprised four steps. We first introduce our study goal with the participants and collect answers to our demographic questions. We then demonstrated our SonifyAR app with an example AR experience. During this phase, participants had the opportunity to ask any questions regarding SonifyAR’s functionalities. Afterwards, participants were asked to independently explore the SonifyAR app. We provided a walking robot model which can be added to indoor surfaces. This model starts walking when tapped. We asked the participants to explore this AR asset freely and use our automatic sound authoring pipeline to create sound effects with the model until they are satisfied with the results. The entire usage process was screen-recorded and the user operations were logged. The final step involved getting participants’ feedback regarding the usability, helpfulness and technical performance of the SonifyAR system based their personal experiences. Participants also provided reasoning and insights to support their answers. The questions and rating results are shown in Figure 7.
@@ 5.3. Results
All participants were able to complete the AR sound authoring task using SonifyAR without difficulty. On average, the authoring process took 406 seconds (SD=137s) to finish. Participants tested an average of 56 (SD=10.48) sound assets, with an average of 6.4 sounds (SD=1.67) assigned to the authoring results.
Feedback on the SonifyAR app was largely positive. All eight participants expressed favorable impressions of the tool. They highly agreed that the SonifyAR tool would be helpful to AR authoring process (Q1,AVG=6 out of 7, SD=0.93). There was also a general willingness for using SonifyAR in their own AR creation practice (Q2, AVG=6.25, SD=1.16). Participants agreed that the sound interaction involving real-world surfaces will improve the immersiveness of AR experiences (Q3, AVG=6.38, SD=1.06). Additionally, they preferred the automated sound authoring process over the manual search for sound assets (Q5, AVG=6.13, SD=0.99). The only area for improvement was the quality of the generated sound (Q4, AVG=4.75, SD=1.39). However, this could be enhanced with the introduction of a more advanced sound generative model, which can be easily integrated into our system. The full list of questions and the respective Likert-type results can be viewed in Figure 7.
Figure 7. Likert-type questions and results of our usability study

### v3

@@ 3.2. Acquisition of AR Sound
While the above design space helps highlight the dimensions of AR sound and current authoring support, below we reflect on key challenges related to the limitations of sound acquisition methods in terms of handling these different dimensions. For example, a mechanical noise of a virtual robot (virtuality) can be easily sourced from local or online sound databases. However, more specific sounds, like a virtual steel ball hitting a physical concrete wall surface or a virtual racecar jumping into a backyard pool (virtuality-reality) requires physical world understanding and material awareness. The space of possible sound effects here is near infinite. Thus, in such cases, generating sounds using text-to-sound models is preferable. Conversely, certain highly recognizable sounds, like a dog barking, pose difficulties for current generative models, often resulting in outputs that are unsatisfactory and noisy. The complexity in user needs and trade-offs in different methods suggests that the combination of both generation and retrieval could serve as a practical solution.
@@ 3.3. Event Representation
To combine generation and retrieval methods, one challenge is how to condition this suite of different sound acquisition methods. While visual formats such as images (Sheffer and Adi, 2023) and videos (Zhou et al., 2018; Ghose and Prevost, 2022) have been widely used as inputs for sound generation models, they fall short in fully capturing AR context information, particularly user actions. To overcome this, we explore the use of text as a universal representation for sound acquisition inputs, which has the following benefits. First, text can sufficiently and precisely convey necessary context information and the specifics of sound-generating events (e.g.,“User slides a ceramic cup on a wooden surface”). Second, given that the AR experience operates within a hardware-software system with multi-modal sensors (e.g., camera, IMU, GPS), we can leverage this system to monitor and log events within the AR experience and summarize them into descriptive text. Third, recent advancements, as demonstrated by works like HuggingGPT (Shen et al., 2023) which uses a LLM to control AI tasks, showcase the practicality of using LLMs as controllers to process text for sound generation.
@@ 4. SonifyAR Implementation
[…]
A pipeline figure showing SonifyAR’s technical process. From left to right: a virtual toy robot on a wooden desk; several smaller figures showing the context information; a text box containing the context; three icons showing the LLM-controlled sound acquisition process, including openAI logo, a search icon, and a gear icon; a simplified UI panel for sound assets selection.
Building on the above design space, we introduce SonifyAR, a novel PbD sound authoring framework that uses context recognition and generative AI to create personalized, context-sensitive sounds for AR interactions (Figure 3). SonifyAR consists of three primary components : (1) Event Textualization: Every user action is recognized as an AR event, and the context of these events is transformed into textual descriptions. (2) Sound Acquisition: An LLM-controlled sound acquisition process that utilize four methods to produces sound effects based on the context. (3) User Interface: An interface that allows users to experience the AR scene and provides the capability to view, modify, and test sound effects for AR events.
@@ 4.1. Event Textualization
In SonifyAR, we utilize a PbD authoring framework to capture potential sound-producing AR events and compile the context information into text. When users interact with the AR space, SonifyAR captures AR events that can lead to sound feedback. Here we define an AR event as a user action and its subsequent result (e.g., when a user taps a virtual object to trigger its animation). For each AR event, we describe it with an event type (e.g., tapping an object ), action source (e.g., user or virtual object), and action target (e.g., virtual object or real-world plane). Information about the involved parties, such as virtual objects or real-world planes, are also included in the context information. This collected event data is then transformed into text with a template.
@@
[…]
Left: a human user holding a phone, while tapping on the screen with another hand. There is a virtual ceramic tea cup on the screen; Right: a text box containing the context information.
@@
To support a broad range of AR sound interactions, we implemented six event types:
@@
    Collide: A virtual object collides with another virtual object or with a real-world surface. We implemented a specialized collision detection mechanism as introduced in subsection 4.4
@@
SonifyAR utilizes ARKit (Apple, 2023a)’s plane detection functionality (Apple, 2023b) to assess the surrounding environment. Since most AR experiences are anchored to planes, the detected plane information serves as a crucial context. To enrich such information, we employ the Deep Material Segmentation model (Upchurch and Niu, 2022) to segment the scene and identify the material of planes (Figure 3), which helps produce realistic sound effects when an AR event involves this plane. Currently, we support six materials: wood, carpet, concrete, paper, metal, and glass. Surfaces that do not fall into one of these six categories are labeled as unknown surface.
@@
Virtual object semantics are also crucial for sound generation. To ensure sound assets align with the virtual object’s material and its state, we collect a text description for all the virtual objects (e.g.,This model is a toy robot made of metal.”) and their animations (e.g.,A toy robot walks.”). These descriptions can be provided by the asset creator or, if necessary, we prompt the user to add relevant details. Any AR event involving the virtual object or its animation will incorporate these text descriptions.
@@
To aggregate the multi-source context information described above, we employ a simple text template that will feed into an LLM-based backend: “This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities].Event Type is the type of event described by the aforementioned type names; Source refers to the subject of the event, which could be the user when the event is directly triggered by user, or virtual object when it interacts with real-world environment; Target Object is the object of the event (e.g., the plane that gets tapped on or the animation that gets played.). Finally, Additional Information on Involved Entities include details that further elucidate the source object, the triggered event, and the target object. Examples include material descriptions and animation details. We leave the corresponding entity field blank when events are missing a target object or additional information.
During user interactions, the system actively monitors and logs events happening in AR space. The context information is fetched and plugged into the text template, crafting a coherent description that reflects the user’s interaction within the AR environment.
@@ 4.2. Sound Acquisition
[…]
A detailed diagram showing SonifyAR’s sound acquisition process. Top shows the context information fed to the LLM; Middle shows a big openAI logo, indicating the LLM; Bottom shows four sound acquisition pipeline and their respective outputs.
@@
SonifyAR utilizes GPT4(OpenAI, 2024) to automatically retrieve or generate context-matching sound assets of an AR event. Inspired by HuggingGPT (Shen et al., 2023) and Visual ChatGPT (Wu et al., 2023b), we utilize the LLM as a controller of multiple sound authoring methods. The LLM takes the text description of the event as input and replies with commands for multiple sound acquisition methods (Figure 5). At our current stage, we support four major sound authoring methods: local recommendation, online retrieval, text2sound generation, and text-guided sound transfer. All sound authoring processes listed below operate concurrently in the backend. This ensures an uninterrupted AR experience.
@@
Local Recommendation. The LLM can recommend sound assets stored in the local database based on semantics in the event description. Similar to other AR authoring tools, we collect a set of sound effects from Adobe Audition’s library(Adobe, 2023b), each labeled with a descriptive filename, like “Crash Aluminum Tray Bang” or “Liquid Mud Suction”. The entire catalog of sound filenames is provided to LLM. When given the event context, the LLM recommends the top five most suitable local sound effects based on their filenames. It then returns the selected sound effect in the format of method1recommend:FILENAME, where FILENAME is replaced by the actual filename. Upon receiving, SonifyAR parses the filename and adds the top corresponding sound as one of the sound options for the event. Users can also long-hold options in the UI to reveal other top recommendation results.
@@
Online Retrieval. We also expand our retrieval capability with an online sound asset database called FreeSound (fre, [n. d.]). We use the FreeSound API, which returns an N-best list of matching sound effects based on a given query. The queries are condensed versions of full event descriptions generated by the LLM, returned in the format method2retrieval:PROMPT, with PROMPT replaced by the specific search query. The returned results are JSON strings. We select the top five matches from the entire set of sound effects returned by the API, which are then downloaded and presented to the user as sound options.
@@
Sound Generation. Beyond recommendation and retrieval methods, we also use the text prompt to generate custom sound effects using an audio diffusion model AudioLDM (Liu et al., 2023). We ask the LLM to compress the event text description into a shortened generation prompt: using the format method3generation:PROMPT, where the PROMPT would be replaced with the generation prompt. Upon receiving such a command from the LLM, SonifyAR sends the prompt to the AudioLDM model, requesting text2sound generation. These newly generated sounds are then shown in the UI as sound options.
@@
Sound-style Transfer. Besides text2audio generation, AudioLDM is also capable of performing text-based sound style transfer. Specifically, for events like tapping, sliding, or colliding, instead of generating a new sound from scratch, SonifyAR uses a default sound effect and initiates a style transfer operation with a text prompt provided by the LLM (e.g., transfer a general tap sound to tapping on glass). This approach allows the output sounds to match the length and rhythm of the input sounds, so that they can be well-timed with actions. Furthermore, when users wish to further modify any provided sound assets, SonifyAR offers text-based style transfer as a fine-tuning and customizing option.
@@ 4.3. User Interface
[…]
Two UI screensots. Left shows a virtual toy plane placed on a real-world wooden desk; Right shows the same interface with a half-transparent call-out panel on the lower half of screen, containing the acquired sound assets for AR events.
As a PbD authoring framework, the SonifyAR application invites users to explore freely with an interactive AR experience (Figure 6 left). Users can interact with the scene, performing actions like moving virtual objects or interacting with the real-world surfaces. When an AR event is detected, a text label appears on the top-left screen to inform the user (Figure 6A) and the sound acquisition component is activated to generate candidate sound effects for the detected events. By clicking the “authoring panel” button (Figure 6D), users can prompt an editing interface (Figure 6F) that overlays the AR scene. The UI includes all detected AR events (Figure 6G) and the corresponding generated sound assets (Figure 6H). Users can click on a sound effect to preview and double click to select and activate it. After confirming their choices, users can hide the authoring panel to resume the AR experience and test the selected sounds. The editing interface can always be re-activated to modify choices.
Additionally, users can long press a sound asset to call out a menu with a suite of exploratory options. For recommended or retrieved sounds, the menu includes an option to list all other sound assets in the top five recommendations. For all sound assets, the menu provides a “style transfer” and a “generate similar sounds” feature, enabling users to style transfer audio effects or to generate similar sounds based on the selected sound. When this feature is selected, users can type a simple text prompt to guide the sound generation process. These choices can be iterated upon or used to explore new sound variations.
@@ 4.4. Technical Implementation
SonifyAR is built with Apple’s ARKit in Swift 5.7 and Xcode 14, and runs as an iOS application on iPhones. Due to the enhanced performance of plane detection on devices with LiDAR (available in iPhone’s Pro lineup starting from iPhone Pro 12), we developed and tested the app on an iPhone 13 Pro Max running iOS 16. We implemented the collision detection between virtual objects and planes with ARKit’s physics simulation (Apple, 2023d). To enable collision detection for animated virtual object parts (e.g. the stomping of a robot foot on the plane), we bind colliders to the joints of the virtual object’s animated bones.
For the LLM-based backend, we use GPT-4.0 (OpenAI, 2024). The text prompt is shown in Appendix A. For material segmentation, we use the Dense Material Segmentation (DMS) model (Upchurch and Niu, 2022) for its detailed material labeling, fast processing time, and high accuracy. For sound generation, we use the AudioLDM model (Liu et al., 2023) due to its state-of-the-art performance and versatility in text2sound generation and text-guided sound2sound transfer. All models are hosted on a server with an NVIDIA RTX 4080 GPU and CUDA 12.2.
@@ 5.1. Participants
We recruited eight participants (six male and two female, aged between 26 and 33) via snowball sampling at Adobe with varying experience in AR. Out of the eight participants, six had prior AR experience, and three had specific experience in AR authoring tools.
@@ 5.2. Procedure
The usability study was conducted in a small meeting room and had four parts: (1) We first introduced our study goal, collected demographic and background information. (2) We then demonstrated SonifyAR with an example AR experience. During this phase, participants had the opportunity to ask any questions regarding SonifyAR’s functionalities. (3) Afterwards, participants were asked to independently explore the SonifyAR app. We provided a walking robot model which could be added to indoor surfaces. This model starts walking when tapped. For consistency, we asked all participants to explore this same AR asset and use our automatic sound authoring pipeline to create sound effects with the model until they were satisfied with the results. The entire usage process was screen-recorded and the user operations were logged. (4) Finally, we sought participant feedback regarding the usability, helpfulness and technical performance of SonifyAR. Participants provided reasoning and insights to support their answers. The questions and rating results are shown in Figure 7.
@@ 5.3. Results
All participants were able to complete the AR sound authoring task using SonifyAR without difficulty. On average, the authoring process took 406 seconds (SD=137s), and the entire study took 35 minutes. Participants tested an average of 56 (SD=10.5) sound assets, with an average of 6.4 sounds (SD=1.7) assigned to the authoring results.
Participant feedback was also largely positive. All eight participants expressed favorable impressions of the tool. They highly agreed that the SonifyAR tool would be helpful to AR authoring process (Q1, avg=6 out of 7, SD=0.93). There was also a general willingness for using SonifyAR in their own AR creation practice (Q2, avg=6.3, SD=1.16). Participants agreed that the sound interaction involving real-world surfaces will improve the immersiveness of AR experiences (Q3, avg=6.4, SD=1.1). Additionally, they preferred the automated sound authoring process over their prior manual search experiences for sound assets (Q5, avg=6.1, SD=1.0). The only area for improvement was the quality of the generated sound (Q4, avg=4.75, SD=1.39). However, this could be enhanced with the introduction of a more advanced sound generative model, which can be easily integrated into our system. The full list of questions and the respective Likert results can be viewed in Figure 7.
Besides the ratings, participants also provided suggestions, especially on the UI and usability. For example, “I really want more usability features that can help communicate these events”, and “I hope there are more associations between the AR events and the sound generation happening in the backend, like some real-time audio hints”.
Figure 7. Likert questions and results of our usability study
[…]
A horizontal chart showing the likert question results with color and numbers.